# Programmable Cognitive Bias in Social Agents

### v2

@@ 9 Limitation
CoBRA represents an important advancement in specifying and regulating LLM-based social agents, empowering the new era of social simulation. We strategically focused on specific aspects. Two key limitations, shaped by this chosen scope, are:
@@
1. Usability: We choose to evaluate CoBRA using technical benchmarks and demonstrations, standard methods for assessing HCI toolkits (Ledo et al., 2018). This paper focuses on exploring the technical feasibility and foundational understanding of the problem. We plan deployment studies with sociologists as the next step to assess CoBRA’s usability (Landay, 2009).
2. Unexplored multi-modal Simulation: We introduce CoBRA for LLM-based agents, where LLMs are utilized for language-based social simulation. We prioritize precision and reproducibility in social simulations. As a result, its applicability to multi-modal human behaviors—such as yawning contagion (observing someone yawn triggers yawning in others)-remains unexplored.
@@
3. Unexplored compositional bias control: We focus on controlling one cognitive bias each time to establish the technical feasibility and precision of CoBRA. As a result, the control of combined biases or interacting constructs (e.g., Authority Effect × Confirmation Bias) remains unexplored. Developing methods to resolve potential conflicts in combined biases is a direction for future research.
4. Partial cognitive bias coverage: This paper introduces CoBRA with a focused set of cognitive biases validated by classic social experiments, ensuring a robust foundation. Expanding to a broader spectrum of biases will further improve CoBRA’s versatility and applicability across diverse social simulation scenarios.
10 Future Work
Building on the advancements presented in this paper, we outline directions for future work to address the above limitations:
1. Utility Assessment through User Studies: We will conduct deployment studies with sociologists and other domain experts. These studies will evaluate CoBRA’s practical value in real-world social simulations, measuring usability, and research impact. By engaging interdisciplinary collaborators, we will iteratively refine CoBRAto improve effectiveness across diverse experimental contexts.
2. Exploration of Multi-Modal Simulation: Future research will integrate vision and audio modalities to simulate non-verbal behaviors, such as yawning contagion, gaze following, and facial-expression-driven emotional contagion. We will develop methodologies for multi-modal agent control and validation protocols to ensure precision and reproducibility in complex, real-world interactions.
3. Compositional Bias Control: We will develop methods for specifying, calibrating, and managing combined biases and interacting constructs. This includes systematic techniques to model and regulate multi-bias compositions (e.g., Authority Effect × Confirmation Bias), alongside conflict-resolution strategies that maintain coherent and realistic agent behavior under competing influences.
4. Expanding Cognitive Bias Dimensions: We will incorporate additional biases such as the Halo Effect, Anchoring Bias, Optimism Bias, and Sunk-Cost Fallacy, and systematically validate them using classic paradigms. This expansion will enhance CoBRA’s versatility and applicability across diverse social simulation applications.
11 Discussion
@@
CoBRA represents an important milestone in the evolution of AI, marking the transition from vague, implicit designs to precise, programmable systems. This section provides a broad discussion of its transformative potential and the ethical considerations critical to ensuring its responsible and impactful use.
Steps Toward a Natural Language Compiler. CoBRA has the potential to introduce a paradigm shift by acting as a natural language compiler, transforming implicit language inputs into precise, reproducible agent behaviors. Traditional LLM-based systems often rely on ambiguous natural language inputs, leading to inconsistent and unpredictable results. CoBRA could bridge this gap, enabling researchers to “compile” natural language into structured and measurable agent specifications. This capability would have far-reaching implications, allowing for scalable agent design that is both interpretable and accessible to non-experts. Future developments might integrate CoBRA with broader NLP pipelines, enabling users to design intelligent agents using plain language instructions while ensuring predictable and consistent outcomes. Such advancements have the potential to revolutionize applications in education, public policy, and training, where realistic and controllable simulations are essential.
Transforming the Role of Cognitive Bias in AI. CoBRA redefines biases as programmable and interpretable features rather than limitations. This reframing enables the creation of socially intelligent agents capable of emulating human-like reasoning and decision-making. For AI, this capability enhances applications such as virtual assistants, interactive simulations, and role-playing systems by making agent behavior more relatable and contextually appropriate. For social science, CoBRA offers a toolkit to study the mechanisms and effects of biases, enabling researchers to simulate complex human behaviors in controlled settings. Ultimately, this transformation positions cognitive biases as tools to achieve deeper insights into both artificial and human intelligence.
@@
Research Ethics. CoBRA’s ability to program biases may raise ethical concerns. Without proper safeguards, these capabilities could be misused to manipulate user behavior, reinforce harmful stereotypes, or amplify societal biases. For instance, agents programmed with biases such as the Bandwagon Effect could be exploited for deceptive marketing or political propaganda. To ensure CoBRA’s responsible use, it is essential to integrate transparency measures, such as bias disclosure, and implement auditing mechanisms to detect and mitigate harmful behaviors. Additionally, ethical guidelines must govern its deployment in sensitive domains, balancing innovation with accountability. At the same time, CoBRA provides an opportunity to model and counteract harmful biases, contributing to the development of fairer and more equitable AI systems. By fostering collaboration between researchers, ethicists, and policymakers, CoBRA can ensure that its capabilities are leveraged for societal benefit while minimizing risks.
12 Conclusion
@@
This paper introduces CoBRA, a novel toolkit that transforms the specification and regulation of LLM-based social agents from implicit natural language descriptions to explicit, programmable control. By treating cognitive biases as programmable features and grounding its approach in validated paradigms from classic social experiments, CoBRA ensures precise, reproducible, and controllable agent behaviors through its Cognitive Bias Index and Behavioral Regulation Engine. Achieving robust performance in technical benchmarks and demonstrations, CoBRA paves the way for practical and impactful applications in social science simulation and beyond.
Appendix A Extended Discussion
This section presents an extended discussion of the background and related works of this paper.
A.1 Cognitive Bias in LLM-As a Feature, Not a Bug
Traditionally, cognitive bias is defined as systematic deviations from rational judgment that lead to suboptimal decision-making outcomes (Unknown, 2015; Baron, 2007). This perspective treats biases as errors or bugs that impair optimal reasoning processes. Recent studies have documented extensive cognitive biases in LLMs, including confirmation bias, anchoring bias, and framing effects (Koo et al., 2023; Boonprakong et al., 2025; Sumita et al., 2025). Research demonstrates that LLMs exhibit amplified cognitive biases in moral decision-making contexts (Cheung et al., 2025), leading to a systematic assessment and mitigation of these biases (Echterhoff et al., 2024).
However, cognitive bias can also be viewed as an adaptive feature that evolved to facilitate social navigation and cooperation (Haselton et al., 2015; Cosmides & Tooby, 1994). From this evolutionary perspective, biases serve important prosocial functions, enabling sophisticated social intelligence rather than representing flaws (Liu et al., 2025b). However, existing research lacks systematic mechanisms to harness these biases as controllable features for enhancing AI social capabilities.
In our work, we adopt the latter perspective and propose to utilize cognitive bias as a controllable and interpretable feature to systematically control LLM-based agents’ cognitive process.
A.2 LLM Behavioral Control Methods
The increasing deployment of LLM-based agents necessitates precise and reliable behavioral control mechanisms. Three prominent techniques for achieving this control are Prompt Engineering, Fine-tuning & Task Vector, and Representation Engineering. Each approach presents a unique set of merits and challenges for guiding model behavior.
Prompt engineering has served as the foundational approach to behavioral control, leveraging natural language instructions to guide model behavior using structured prompting strategies (Rath et al., 2025; Zhuo et al., 2024). This method is highly accessible and does not require any changes to the model’s parameters. However, designing effective prompts has many pitfalls (Zamfirescu-Pereira et al., 2023), and fundamental limitations persist: LLMs exhibit extreme sensitivity to prompt variations, with large accuracy differences from minor changes (Sclar et al., 2024).
Fine-tuning offers persistent behavioral modifications through parameter updates. Techniques such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have become dominant for model alignment (Ouyang et al., 2022; Rafailov et al., 2023). Task vectors enable compositional behavioral modifications through arithmetic operations on weight differences (Ilharco et al., 2023). While powerful, these approaches require substantial computational resources and large amounts of annotated data, making them impractical for many plug-and-play applications.
Representation engineering provides a method for behavioral control through direct manipulation of a model’s internal representations during inference (Zou et al., 2023). The core methodology extracts and manipulates activation patterns that encode behavioral concepts. Activation steering methods compute steering vectors from contrasting behavioral examples (Rimsky et al., 2024), which can be applied during inference to modulate behavior without parameter updates (Todd et al., 2024). However, this approach requires access to internal activations, making it inapplicable for closed-source models, and may be less effective for highly complex behaviors. Recent applications include personality control (Chen et al., 2025), community-specific preference modeling (Li et al., 2025), harmlessness control (Arditi et al., 2024), and multi-objective alignment (Shi et al., 2024). These methods operate with minimal inference overhead while maintaining base model capabilities.
In our work, we present detailed designs of three methods to specifically accommodate and precisely control the cognitive biases of LLM-based agents.
A.3 LLM-based Social Simulation & Agent Specification
Existing work on LLM-based agent social simulations typically instantiates LLMs as autonomous agents by assigning them scenario-specific prompts, roles, and objectives (Park et al., 2023; Chen et al., 2024b). Methodologically, these studies typically follow a pipeline consisting of three core stages: 1. Agent Specification, 2. Social Environment Set-up, and 3. Behavior Observation. During Agent Specification, researchers harness the expressiveness and flexibility of natural language to define agent identity, knowledge, and behavioral tendencies, enabling the creation of contextualized personas. In the Social Environment Set-up stage, agents are embedded in experimental scenarios inspired by social science literature (Liu et al., 2025b), or in constructed social environments with specific rules and resources (Zhou et al., 2024; Park et al., 2023; Chen et al., 2024b), such as sandbox simulations or interactive narratives. Finally, in the Behavior Observation stage, researchers examine emergent social dynamics (Zhu et al., 2025)—including cooperation, norm formation, conflict, and human-like heuristic behaviors such as conformity, by observing interactions both among agents and between agents and human participants.
Within this pipeline, Agent Specification is crucial for defining and shaping agent behavior, representative examples include: persona-based specification (Rath et al., 2025; Park et al., 2023; Zhou et al., 2024) and role-based specification (Li et al., 2023a; Zheng et al., 2023; Chen et al., 2024b; Jia et al., 2024). Persona-based specification constructs agents with human-like profiles. Core elements typically include demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five or MBTI), personal interests (hobbies, preferences), and social relationships (friends, family, colleagues) that situate the agent in a social context. In practice, persona detail can range from a single facet (such as demographics) to comprehensive architectures—such as Generative Agents (Park et al., 2023)—that integrate multiple attributes with memory to produce situated behavior. Role-based specification defines agents primarily by their professional or task-oriented identities, rather than rich personal backgrounds (Chen et al., 2024b). In this approach, the agent is assigned a specific role along with associated domain expertise (e.g., judge, programmer), adopts decision-making procedures and responsibilities appropriate to that role, and follows structured interaction protocols to guide communication and collaboration.
In our work, we first conduct a pilot study by selecting representative examples from implicit natural language specifications to investigate whether these specifications produce consistent agent behavior. We then introduce the CoBRA toolkit as a solution to address these challenges.
Appendix B Pilot Study Detail
@@
B.1 Example of Persona-Based Specification
@@
B.2 Example of Role-Based Specification
”Teacher” agent specification (Chen et al., 2024b):
You are an experienced dialogue teacher. As a good teacher, you carefully check the correctness of the given response based on the text. When the solution is flawed, you should patiently teach the students how to give better response. You must respond in the following format:
Interesting: (a score between 0 and 9)
Engaging: (a score between 0 and 9)
Specific: (a score between 0 and 9)
Relevant: (a score between 0 and 9)
Semantically Appropriate: (a score between 0 and 9)
Understandable: (a score between 0 and 9)
Fluent: (a score between 0 and 9)
Overall Impression: (a score between 0 and 9)
Advice: (your advice on how to correct the solution
Problem: \$task_description
Student’s Solution: \$solution
Now carefully check the student’s solution, and give your response.
Appendix C Classic Social Experiment Testbed Detail
@@
C.1 Bandwagon Effect Testbed
@@ 1. Asch’s Line Paradigm: Testing Bandwagon Effect via Group Consensus
We construct 15 distinct [Statement] items across three thematic categories—Future, Open-Ended, and Common Sense—mirroring the design used in the Authority Effect Testbed, resulting in a total of 15 unique prompt variations.
@@
$\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (12)
@@ 2. Hotel Towel Paradigm: Testing Bandwagon Effect via Majority Behavior
$\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{10}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (13)
C.2 Confirmation Bias Testbed

### v3

@@ 9. LIMITATIONS & FUTURE WORK
CoBRA represents an important advancement in specifying and regulating LLM-based social agents, empowering the new era of social simulation. We strategically focused on specific aspects. Key limitations of this scope are:
@@
Usability. We choose to evaluate CoBRA using technical benchmarks and demonstrations, standard methods for assessing HCI toolkits (Ledo et al., 2018). This paper focuses on exploring the technical feasibility and foundational understanding of the problem. We plan deployment studies with sociologists as the next step to assess CoBRA’s usability (Landay, 2009).
Unexplored multi-modal Simulation. We introduce CoBRA for LLM-based agents, where LLMs are utilized for language-based social simulation. We prioritize precision and reproducibility in social simulations. As a result, limits the social phenomena we can model (e.g., non-verbal contagion, gaze following). Future work will extend CoBRA to multi-modal settings (vision and audio), with corresponding control and evaluation methods.
@@
Unexplored compositional bias control. We focus on controlling one cognitive bias at a time to establish the technical feasibility and precision of CoBRA, so combined biases and interacting constructs (e.g., Authority Effect × Confirmation Bias) remain unexplored. Future work will develop compositional control methods that allow multiple interacting biases with principled conflict-resolution strategies.
Model accessibility. While CoBRA supports behavioral control across input, activation, and parameter spaces, the latter two require access to model internals and are therefore limited to open-source LLMs. Thus, closed-source models are limited to input-space (prompt-based) control, with reduced precision.
10. Discussion
@@
CoBRA represents an important milestone in the evolution of AI, marking the transition from vague, implicit designs to precise, programmable systems. This section provides a broad discussion of its methodological position, transformative potential, and ethical considerations critical to ensuring its responsible use.
CoBRA as a Behavioral Control Layer. CoBRA controls how strongly contextual cues (framing, authority, majority, etc.) shape an agent’s behavior; it does not specify underlying cognitive mechanisms. For example, in cases such as the Stanford Prison Experiment (Haslam et al., 2019), CoBRA can be used to control an agent’s susceptibility to authority or majority influence in a quantitative and reproducible way, but it does not by itself adjudicate between competing psychological explanations (e.g., role-based accounts vs. identity leadership accounts) of why such contextual factors produce harm in humans. Our Cognitive Bias Index (CBI), therefore, captures changes in observable behavioral tendencies (e.g., how often an agent selects options aligned with an authority cue), rather than pinpointing which specific internal “mechanism” in the LLM has been altered. For example, when an authority-related CBI increases, this may reflect some combination of stronger role endorsement, greater deference to explicitly stated instructions, or other learned representations; CoBRA does not distinguish among these possibilities.
CoBRA should therefore be seen as a behavioral control layer for LLM-based agents: it makes context sensitivity programmable, but claims about human mental processes must ultimately be grounded in human data and theory. Social simulations using CoBRA can generate hypotheses and stress-test theories under varied settings, but they do not by themselves constitute evidence about human mechanisms.
Steps Toward Agent Specification Compiler. Our motivation for building such a control layer is that current LLMs still exhibit limited and unstable perceptual capabilities; without explicit controls, experimenters lack quantitative levers over agent specifications across conditions, which weakens the rigor of social simulations. In this sense, CoBRA can be viewed as a lightweight Agent Specification Compiler: it helps transform high-level natural-language descriptions of cognitive tendencies into structured, reproducible behavioral specifications via CBIs.
The current performance and accuracy of CoBRA are not final. Incorporating additional classic experimental paradigms, systematically augmenting prompts, and combining CoBRA with complementary control methods are promising directions to further improve the fidelity, robustness, and coverage of behavioral regulation.
Research Ethics. CoBRA’s ability to program biases may raise ethical concerns. Without proper safeguards, it could be misused to manipulate user behavior, reinforce harmful stereotypes, or amplify societal biases. For example, agents programmed with the Bandwagon Effect could be exploited for deceptive marketing or political propaganda. Similar risks arise in research practice: because CoBRA allows systematic control over susceptibility to contextual cues, CBI could be misused to “dial in” desired outcomes (e.g., by tuning CBI post hoc to confirm prior expectations). To mitigate this, CBI settings should be treated as explicit experimental parameters—pre-registered where possible, systematically varied, and transparently reported—rather than adjusted after the fact to fit a preferred narrative.
@@
To ensure CoBRA’s responsible use, it is important to adopt transparency measures (e.g., bias disclosure) and auditing mechanisms to detect and mitigate harmful behaviors, and to develop ethical guidelines for deployment in sensitive domains. At the same time, CoBRA provides an opportunity to model and counteract harmful biases, supporting the development of fairer and more equitable AI systems. By fostering collaboration between researchers, ethicists, and policymakers, CoBRA’s capabilities can be leveraged for societal benefit while minimizing risks.
11. Conclusion
@@
This paper introduces CoBRA, a novel toolkit that transforms the specification and regulation of LLM-based social agents from implicit natural language descriptions to explicit, programmable control. At the heart of CoBRA is a novel closed-loop system primitive that continuously measures the amount of cognitive biases demonstrated in agents’ responses to a set of validated classic social science experiments and adjusts the agent specifications accordingly. CoBRA ensures precise, reproducible, and controllable agent behaviors through its Cognitive Bias Index and Behavioral Regulation Engine. Achieving robust performance in technical benchmarks and demonstrations, CoBRA paves the way for practical and impactful applications in social science simulation and beyond.
Appendix A Pilot Study Detail
@@
A.1. Example of Persona-Based Specification
@@
A.2. Statistical Analysis Details
A.2.1. Weighted frequency difference Computation
The weighted frequency difference (WFD) quantifies preference bias on a continuous scale:
(7) $\mathrm{WFD}$ $=2f(\text{Strongly}_{A})+f(\text{Somewhat}_{A})$
$\quad-f(\text{Somewhat}_{B})-2f(\text{Strongly}_{B})$
where $f(\cdot)$ represents the proportion of responses in each category, and $A$ and $B$ denote the two framing conditions. The score ranges from 2-2 (complete preference for $B$) to +2+2 (complete preference for $A$), with 0 indicating no bias.
A.2.2. Hypothesis Testing Framework
  • Hypothesis 1: Profile Effect on Bias. We tested whether profiles produce meaningful effects using TOST (Two One-Sided Tests) equivalence testing on weighted frequency difference. We decompose this into two checks:
    • H1a: Behavioral Distinctness (Economist vs Blank). We test if the Economist profile is statistically different from the Blank baseline and reduce bias compared to the Blank baseline.
    • H1b: Reduced Susceptibility (Economist vs Common). We test if the Economist profile is statistically different from the Common profile and reduce bias compared to the Common baseline.
    For both hypotheses, we use the following bounds:
    • $H_{0}$: $|\mu_{profile}-\mu_{baseline}|\geq\Delta$ (Meaningful difference exists)
    • $H_{1}$: $|\mu_{profile}-\mu_{baseline}|<\Delta$ (Effect is within equivalence bounds, i.e., no meaningful difference)
    Effect size threshold: $\Delta=0.5$ scale units (corresponding Cohen’s $d\approx 0.33$, pooled $\sigma=1.498$). Sample size: $n=150$ per condition. Significance: $\alpha=0.05$.
  • Hypothesis 2: Cross-Model Consistency. Chi-square tests on response distributions (4 models × 5 choice categories):
    • $H_{0}$: Response patterns independent of model
    • $H_{1}$: Patterns differ across models
Model Profile NB Baseline NB Cohen’s $d$ $p_{equiv}$ Verdict
H1a: Economist vs Blank (TOST equivalence testing with $\Delta=0.5$ Likert units)
Mistral-7B 63.33 68.00 0.03 0.004 EQUIVALENT
Gemma-2-9B 27.33 19.33 -0.05 0.007 EQUIVALENT
DeepSeek-v3 40.00 56.67 0.11 0.024 EQUIVALENT
GPT-4o-mini 28.67 60.00 0.20 0.148 Not PASS
Summary 3/4 Equivalent 1/4 Distinct
H1b: Economist vs Common (TOST equivalence testing with $\Delta=0.5$ Likert units)
Mistral-7B 63.33 59.33 -0.03 0.003 EQUIVALENT
Gemma-2-9B 27.33 41.33 0.10 0.017 EQUIVALENT
DeepSeek-v3 40.00 34.00 -0.04 0.006 EQUIVALENT
GPT-4o-mini 28.67 62.00 0.22 0.172 Not PASS
Summary 3/4 Equivalent 1/4 Distinct
Cross-Model Consistency (Chi-square test on response distributions)
Economist profile Chi-square $\chi^{2}=30.96$ $\mathbf{p=0.002}$ INCONSISTENT
Common profile Chi-square $\chi^{2}=26.32$ $\mathbf{p=0.010}$ INCONSISTENT
Blank profile Chi-square $\chi^{2}=43.82$ $\mathbf{p<0.001}$ HIGHLY INCONSISTENT
Table 4. Statistical test results using TOST equivalence testing. Cohen’s $d$ measures standardized effect size; $p_{equiv}$ reports TOST equivalence test results (max of two one-sided p-values). H1a (Eco vs Blank): Three models (Mistral, Gemma, DeepSeek) show statistically equivalent effects to Baseline ($p_{equiv}<0.05$), indicating the persona failed to produce distinct behavior. H1b (Eco vs Common): The same three models show equivalence to the Common profile, indicating the Economist persona did not reduce susceptibility compared to a layperson. Only GPT-4o-mini consistently showed non-equivalent (distinct) behavior.
A.2.3. Interpretation Guidelines
  • TOST framework: Two one-sided tests determine if the effect falls within equivalence bounds $\Delta,+\Delta]$. A significant result ($p<0.05$) indicates Equivalence (i.e., the Economist persona had no meaningful effect of reducing bias).
  • Key findings: Implicit specification did not reliably yield expected behaviors across models.
    • GPT-4o-mini: The only model to pass heuristic checks.
    • Mistral-7B: Failed to differentiate. Economist, Common, and Blank agents behaved quite similarly, with all profiles statistically equivalent to the baseline ($p_{equiv}<0.01$).
    • DeepSeek-v3: Failed susceptibility check. Economist agents were more susceptible to framing effects (Net Bias 40.00) than Common agents (Net Bias 34.00), contradicting domain expertise expectations.
    • Gemma-2-9B: Failed distinctness check. Economist behavior was closer to Blank agents (d=0.05d=-0.05) than to Common agents ($d=0.10$), failing to exhibit the expected expert deviation.
Appendix B Classic Social Experiment Testbed Detail
@@
Definition 0 (Cognitive Bias Index).
(8) $\left\{\mathrm{CBI}_{b}^{(k)}\;\middle|\;b\in\mathcal{B},\ k=1,2,\ldots,m\right\}$
(9) $\mathrm{CBI}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
B.1. Bandwagon Effect Testbed
@@ 1. Asch’s Line Paradigm: Testing Bandwagon Effect via Group Consensus
We construct 15 distinct [Statement] items across three thematic categories: Future, Open-Ended, and Common Sense, mirroring the design used in the Authority Effect Testbed, resulting in a total of 15 unique prompt variations.
@@
(10) $\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
@@ 2. Hotel Towel Paradigm: Testing Bandwagon Effect via Majority Behavior
(11) $\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{10}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
B.2. Confirmation Bias Testbed